suppose the weight of 800 students are normally distributed with mean 75kg and standard deviation 5kg. find the numer of students that weigh more than 80kg


1
Expert's answer
2022-01-25T16:46:41-0500

Let "X=" the student's weight: "X\sim N(\mu, \sigma^2)."

Given "\mu=75\ kg, \sigma=5\ kg."


"P(X>80)=1-P(X\leq 80)"

"=1-P(Z\leq\dfrac{80-75}{5} )=1-P(Z\leq 1)"

"\approx0.158655"


"800(0.158655)=127"

127 students weigh more than 80kg.